Group Selection Page
HELP
Controls Group
ITDM
Vacuum
WP78
FXE
SPB
SQS
SXP
XO
Migrated2Zulip
Archived2GitlabPages
Archived2PDF
not2MigrateOrArchive
MigratedByGroup
SCS
SXP Migrated2Zulip
Photon Commissioning Team
XRO
DOC
HED Migrated2Zulip
WP74
MID
PSPO_XO
XO_Operation
SQS Migrated2Zulip
FXE Migrated2Zulip
DET
Sample Environment
SPB_mig
SASE1 Status and SASE3 Readiness
Technical integration and coordination
Joint Operation and Readiness Logbook, Contact: T.Haas, A. Violante, G. Wellenreuther
Not logged in
List
|
Find
|
Login
|
Help
Message ID:
529
Entry time:
30 May 2018, 10:55
Author:
Adriano Violante
Location:
General
Severity:
Information
Subject:
Status report after dispatch meeting 09.30
AE
CRLs in XTD1 being implemented
GATT and EPS loop in SASE3 being updated
SCS
Tests continue - No issues so far
SQS
ELMO controllers have to be configured for the mirror motors test. Slit motors have lower priority. Cables can be tested w/o being physically connected to the motors, but to configure the ELMO controllers the cables have to be connected to the real motors. Tomorrow AE is available to support these activities.
Basler cameras set up and working.
ITDM
SPB AGIPD test today or tomorrow
CAS
Hot fix for SPB released yesterday. This will address the issue of crushings when instantiating a new device. Has to be tested with SPB.
XTD 6 will have no electricity for the racks today since Wille is completing the cabling. MTCA being shut down accordingly.
ELOG V3.1.4-7c3fd00